# MT-Depth: Multi-task Instance feature analysis for the Depth Completion

### v1

@@ Abstract
Depth completion plays a vital role in 3D perception systems, especially in scenarios where sparse depth data must be densified for tasks such as autonomous driving, robotics, and augmented reality. While many existing approaches rely on semantic segmentation to guide depth completion, they often overlook the benefits of object-level understanding. In this work, we introduce an instance-aware depth completion framework that explicitly integrates binary instance masks as spatial priors to refine depth predictions. Our model combines four main components: a frozen YOLO V11 instance segmentation branch, a U-Net-based depth completion backbone, a cross-attention fusion module, and an attention-guided prediction head. The instance segmentation branch generates per-image foreground masks that guide the depth branch via cross-attention, allowing the network to focus on object-centric regions during refinement. We validate our method on the Virtual KITTI 2 dataset, showing that it achieves lower RMSE compared to both a U-Net-only baseline and previous semantic-guided methods, while maintaining competitive MAE. Qualitative and quantitative results demonstrate that the proposed model effectively enhances depth accuracy near object boundaries, occlusions, and thin structures. Our findings suggest that incorporating instance-aware cues offers a promising direction for improving depth completion without relying on dense semantic labels.
@@ 1 Introduction
Many CNN-based depth completion models adopt an encoder–decoder architecture that takes RGB and sparse depth as input, using skip connections to preserve spatial details klingner2020selfsupervised; Eldesokey2020confidence; Li2024SemanticGuidedDC. Despite their success, single-task models often struggle in challenging scenarios—such as thin structures, distant objects, and occlusion boundaries—where geometric cues alone are insufficient. To overcome these challenges, recent research has explored the integration of semantic information to guide depth estimation AtapourAbarghouei2019ToCO; Huang2019indoor. Semantic segmentation introduces class-level context, enabling the network to recognize object boundaries and propagate meaningful structure across regions of missing data.
Multi-task learning (MTL) has emerged as a natural extension, allowing models to jointly learn depth, segmentation, and other geometric cues within a shared framework ZHAO2023106496; Moeskops2016; Shivakumar2019; Cheng2019CSPNLC. By leveraging shared representations across related tasks, MTL not only improves generalization but also facilitates information transfer between domains. For instance, predicting surface normals alongside depth can enhance 3D consistency, while semantic segmentation helps identify meaningful object boundaries that align with depth discontinuities.
Among these, SemSegDepth Lagos_2022 is a notable example that jointly predicts dense depth and semantic segmentation maps from RGB and sparse depth inputs. It uses a shared encoder with separate decoder branches for each task and demonstrates that fusing semantic features improves depth quality. Similarly, PanDepth Lagos2022PanDepthJP extends the idea to panoptic segmentation, integrating semantic, instance, and depth understanding into a single multi-task model.
@@
In real-world applications like robotics and autonomous driving, instance-level understanding is critical. Accurate depth prediction around objects—especially for small, distant, or partially occluded entities—is essential for collision avoidance, manipulation, and interaction. Therefore, combining instance segmentation with depth completion offers the potential to significantly enhance 3D scene understanding Lee2021LearningMD; Jo2024; s24072374.
Recent advances in real-time instance segmentation, particularly the YOLO family, provide efficient and accurate object detection and mask prediction Bochkovskiy2020yolov4; khanam2024yolov5deeplookinternal; YOLOv8. YOLO V11, a recent version, introduces instance mask heads alongside object detection outputs, enabling fast extraction of binary masks at multiple scales Khanam2024YOLOv11AO. These instance-aware features are well-suited for integration with depth completion, as they provide per-object guidance without requiring pixel-level semantic class labels.
@@
We validate our model on the Virtual KITTI 2 dataset cabon2020vkitti2; Gaidon2016vkitti, a synthetic benchmark offering perfect ground truth for depth, instance segmentation, and RGB under varied lighting and weather conditions. Our results show that the proposed instance-aware architecture outperforms baseline depth completion models and achieves competitive accuracy compared to semantic-guided approaches—without requiring class-level labels. Our model reduces RMSE while maintaining low MAE, especially in regions near object edges and thin structures.
@@ 2.1 Instance Segmentation
Recent progress in instance segmentation has accelerated thanks to advances in architectures, loss functions, and data-efficient learning strategies. A 2025 survey reviews over 200 deep‑learning based segmentation models, categorizing them by backbone, mask decoding, and temporal consistency strategies — a useful resource to track current trends and limitations of mask-based segmentation methods. XIA2025128740; tomography11050052 At the same time, novel frameworks like diffusion‑based instance segmentation have emerged: the authors of a 2025 study demonstrate that diffusion models can generate high‑quality instance masks even for challenging cases such as overlapping objects or weak boundaries. Ma2025diffusion For transparent or translucent objects — traditionally hard to segment — a recent few‑shot method leverages data augmentation and template consistency to reliably produce instance masks from limited examples. Cherian2024TrInSeg On the application side, instance segmentation remains fundamental for tasks such as robot bin picking, scene understanding, and autonomous perception. Xu2025VIS Finally, contemporary surveys of segmentation methods highlight the trade‑offs among speed, mask fidelity, and generalization. Wolk2024survey
@@ 2.2 Depth Completion
Depth completion remains a hot topic as it bridges sparse sensor outputs (e.g. LiDAR or partial depth) with dense depth estimation required for perception or 3D reconstruction. A newly proposed method from late 2024, SigNet, treats depth completion as a two‑step problem: first densify sparse depth using non‑CNN tools, then enhance the coarse depth via a CNN. This “degradation-aware” paradigm helps alleviate artifacts caused by direct CNN completion from very sparse inputs. Yan2024SigNet Meanwhile, hybrid depth‑guided sensor fusion approaches are also gaining traction, this approach describes a model that combines LiDAR and camera data with semantic and geometric fusion to achieve robust perception under challenging conditions. Tim2025 On the 3D completion side, a recent method integrates stereo depth cues and implicit depth-aware features to complete full 3D occupancy and semantic maps — a promising direction for applications like autonomous driving and scene reconstruction. Miao2023OccDepth Additionally, transformer-based and multi‑path aggregation networks have been developed for semantic scene completion, demonstrating strong performance on large-scale driving datasets. Xia2023SCPNet Finally, a 2023 survey of depth and scene reconstruction methods outlines major challenges and categorizes modern approaches, providing useful context for evaluating new models. Rajapaksha2024
@@ 2.3 Multi‑Task Learning
Joint learning of multiple related tasks — such as depth estimation, segmentation, surface normals, or scene reconstruction — has shown consistent benefits by enabling cross-task information sharing and improved generalization. A 2025 work proposes a novel multi-task framework with a shared encoder and separate decoders for surface normals and depth, demonstrating improved consistency across depth and surface predictions CHAVARRIASSOLANO2025103379. Meanwhile, the transformer-based framework SwinMTL (2024) performs concurrent depth estimation and semantic segmentation on monocular images, using adversarial training to refine predictions and showing strong performance across both tasks taghavi2024swinmtl. Combining depth completion and object detection has also been explored: a 2024 paper introduces a foreground‑aware network that fuses depth and object detection tasks using a transformer-based architecture, improving both depth completion and detection under sparse data conditions. Pan2024ForegroundAware In addition, generative‑adversarial methods have been proposed to jointly learn segmentation and depth from monocular images, showing that adversarial signals can improve the realism and consistency of depth maps. Zhang2021MTGAN Finally, a broader survey of multi‑task dense prediction methods published in 2024 outlines key challenges in balancing losses, architecture design, and cross-task interference. li2024multitask
@@ 3.1 Segmentation Branch Using YOLO V11
YOLO (You Only Look Once) is a real-time object detection framework that simultaneously performs object classification and localization in a single pass through a neural network Khanam2024YOLOv11AO. It divides the input image into a fixed grid and predicts bounding boxes, objectness scores, and class probabilities directly for each cell. This one-stage design contrasts with earlier two-stage detectors like R-CNN, making YOLO significantly faster while maintaining strong accuracy. Over time, the YOLO family has evolved through various iterations (v1 to v10), each improving on backbone architecture, feature fusion, and multi-scale detection performance.
@@ 3.2 Depth Completion Branch Using U-Net
The depth completion branch is responsible for estimating dense depth maps from sparse depth inputs and RGB images Ronneberger2015; Wang2022; Ramesh2023SIUNetSI. This is achieved using a modified U-Net architecture, which takes as input a concatenation of the RGB image, sparse depth, and a binary mask indicating valid depth pixels. Let $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ denote the RGB image, $\mathbf{D}_{\text{sparse}}\in\mathbb{R}^{1\times H\times W}$ the sparse depth map, and $\mathbf{V}\in\{0,1\}^{1\times H\times W}$ the corresponding validity mask. The combined input to the U-Net is:
@@ 4 Results and Experiment
4.1 Kitti Virtual 2 Dataset
We use the Virtual KITTI 2 dataset, a photorealistic synthetic clone of real-world KITTI scenes, designed for training and evaluating vision models under controlled conditions cabon2020vkitti2; Gaidon2016vkitti. It provides synchronized RGB images, dense depth maps, instance segmentation, and class-level semantic labels across multiple virtual scenes and weather conditions. Each frame is rendered from multiple cameras and includes perfect ground truth, making it ideal for depth completion tasks. In our experiments, we use data from the front-facing camera (Camera_0), resizing all inputs to $256\times 512$ and applying an 80/20 split for training and validation 1.

### v2

@@ Abstract
Depth completion plays a vital role in 3D perception systems, especially in scenarios where sparse depth data must be densified for tasks such as autonomous driving, robotics, and augmented reality. While many existing approaches rely on semantic segmentation to guide depth completion, they often overlook the benefits of object-level understanding. In this work, we introduce an instance-aware depth completion framework that explicitly integrates binary instance masks as spatial priors to refine depth predictions. Our model combines four main components: a frozen YOLO V11 instance segmentation branch, a U-Net-based depth completion backbone, a cross-attention fusion module, and an attention-guided prediction head. The instance segmentation branch generates per-image foreground masks that guide the depth branch via cross-attention, allowing the network to focus on object-centric regions during refinement. We validate our method on the Virtual KITTI 2 dataset, showing that it achieves lower Root Mean Squared Error (RMSE) compared to both a U-Net-only baseline and previous semantic-guided methods, while maintaining competitive Mean Absolute Error (MAE). Qualitative and quantitative results demonstrate that the proposed model effectively enhances depth accuracy near object boundaries, occlusions, and thin structures. Our findings suggest that incorporating instance-aware cues offers a promising direction for improving depth completion without relying on dense semantic labels.
@@ 1 Introduction
Many CNN-based depth completion models adopt an encoder–decoder architecture that takes RGB and sparse depth as input, using skip connections to preserve spatial details [1, 2, 3]. Despite their success, single-task models often struggle in challenging scenarios—such as thin structures, distant objects, and occlusion boundaries—where geometric cues alone are insufficient. To overcome these challenges, recent research has explored the integration of semantic information to guide depth estimation [4, 5]. Semantic segmentation introduces class-level context, enabling the network to recognize object boundaries and propagate meaningful structure across regions of missing data.
Multi-task learning (MTL) has emerged as a natural extension, allowing models to jointly learn depth, segmentation, and other geometric cues within a shared framework [6, 7, 8, 9]. By leveraging shared representations across related tasks, MTL not only improves generalization but also facilitates information transfer between domains. For instance, predicting surface normals alongside depth can enhance 3D consistency, while semantic segmentation helps identify meaningful object boundaries that align with depth discontinuities.
Among these, SemSegDepth [10] is a notable example that jointly predicts dense depth and semantic segmentation maps from RGB and sparse depth inputs. It uses a shared encoder with separate decoder branches for each task and demonstrates that fusing semantic features improves depth quality. Similarly, PanDepth [11] extends the idea to panoptic segmentation, integrating semantic, instance, and depth understanding into a single multi-task model.
@@
In real-world applications like robotics and autonomous driving, instance-level understanding is critical. Accurate depth prediction around objects—especially for small, distant, or partially occluded entities—is essential for collision avoidance, manipulation, and interaction. Therefore, combining instance segmentation with depth completion offers the potential to significantly enhance 3D scene understanding [12, 13, 14].
Recent advances in real-time instance segmentation, particularly the YOLO family, provide efficient and accurate object detection and mask prediction [15, 16, 17]. YOLO V11, a recent version, introduces instance mask heads alongside object detection outputs, enabling fast extraction of binary masks at multiple scales [18]. These instance-aware features are well-suited for integration with depth completion, as they provide per-object guidance without requiring pixel-level semantic class labels.
@@
We validate our model on the Virtual KITTI 2 dataset [19, 20], a synthetic benchmark offering perfect ground truth for depth, instance segmentation, and RGB under varied lighting and weather conditions. Our results show that the proposed instance-aware architecture outperforms baseline depth completion models and achieves competitive accuracy compared to semantic-guided approaches—without requiring class-level labels. Our model reduces RMSE while maintaining low MAE, especially in regions near object edges and thin structures.
@@ 2.1 Instance Segmentation
Recent progress in instance segmentation has accelerated thanks to advances in architectures, loss functions, and data-efficient learning strategies. A 2025 survey reviews over 200 deep‑learning based segmentation models, categorizing them by backbone, mask decoding, and temporal consistency strategies — a useful resource to track current trends and limitations of mask-based segmentation methods. [21, 22] At the same time, novel frameworks like diffusion‑based instance segmentation have emerged: the authors of a 2025 study demonstrate that diffusion models can generate high‑quality instance masks even for challenging cases such as overlapping objects or weak boundaries. [23] For transparent or translucent objects — traditionally hard to segment — a recent few‑shot method leverages data augmentation and template consistency to reliably produce instance masks from limited examples. [24] On the application side, instance segmentation remains fundamental for tasks such as robot bin picking, scene understanding, and autonomous perception. [25] Finally, contemporary surveys of segmentation methods highlight the trade‑offs among speed, mask fidelity, and generalization. [26]
@@ 2.2 Depth Completion
Depth completion remains a hot topic as it bridges sparse sensor outputs (e.g. LiDAR or partial depth) with dense depth estimation required for perception or 3D reconstruction. A newly proposed method from late 2024, SigNet, treats depth completion as a two‑step problem: first densify sparse depth using non‑CNN tools, then enhance the coarse depth via a CNN. This “degradation-aware” paradigm helps alleviate artifacts caused by direct CNN completion from very sparse inputs. [27] Meanwhile, hybrid depth‑guided sensor fusion approaches are also gaining traction, this approach describes a model that combines LiDAR and camera data with semantic and geometric fusion to achieve robust perception under challenging conditions. [28] On the 3D completion side, a recent method integrates stereo depth cues and implicit depth-aware features to complete full 3D occupancy and semantic maps — a promising direction for applications like autonomous driving and scene reconstruction. [29] Additionally, transformer-based and multi‑path aggregation networks have been developed for semantic scene completion, demonstrating strong performance on large-scale driving datasets. [30] Finally, a 2023 survey of depth and scene reconstruction methods outlines major challenges and categorizes modern approaches, providing useful context for evaluating new models. [31]
@@ 2.3 Multi‑Task Learning
Joint learning of multiple related tasks — such as depth estimation, segmentation, surface normals, or scene reconstruction — has shown consistent benefits by enabling cross-task information sharing and improved generalization. A 2025 work proposes a novel multi-task framework with a shared encoder and separate decoders for surface normals and depth, demonstrating improved consistency across depth and surface predictions [32]. Meanwhile, the transformer-based framework SwinMTL (2024) performs concurrent depth estimation and semantic segmentation on monocular images, using adversarial training to refine predictions and showing strong performance across both tasks [33]. Combining depth completion and object detection has also been explored in [34], this work introduces a foreground‑aware network that fuses depth and object detection tasks using a transformer-based architecture, improving both depth completion and detection under sparse data conditions. In addition, generative‑adversarial methods have been proposed to jointly learn segmentation and depth from monocular images, showing that adversarial signals can improve the realism and consistency of depth maps. [35] Finally, a broader survey of multi‑task dense prediction methods published in 2024 outlines key challenges in balancing losses, architecture design, and cross-task interference. [36]
@@ 3.1 Segmentation Branch Using YOLO V11
YOLO (You Only Look Once) is a real-time object detection framework that simultaneously performs object classification and localization in a single pass through a neural network [18]. It divides the input image into a fixed grid and predicts bounding boxes, objectness scores, and class probabilities directly for each cell. This one-stage design contrasts with earlier two-stage detectors like R-CNN, making YOLO significantly faster while maintaining strong accuracy. Over time, the YOLO family has evolved through various iterations (v1 to v10), each improving on backbone architecture, feature fusion, and multi-scale detection performance.
@@ 3.2 Depth Completion Branch Using U-Net
The depth completion branch is responsible for estimating dense depth maps from sparse depth inputs and RGB images [37, 38, 39]. This is achieved using a modified U-Net architecture, which takes as input a concatenation of the RGB image, sparse depth, and a binary mask indicating valid depth pixels. Let $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ denote the RGB image, $\mathbf{D}_{\text{sparse}}\in\mathbb{R}^{1\times H\times W}$ the sparse depth map, and $\mathbf{V}\in\{0,1\}^{1\times H\times W}$ the corresponding validity mask. The combined input to the U-Net is:
@@ 4 Results and Experiment
4.1 Virtual KITTI 2 Dataset
We use the Virtual KITTI 2 dataset, a photorealistic synthetic clone of real-world KITTI scenes, designed for training and evaluating vision models under controlled conditions [19, 20]. It provides synchronized RGB images, dense depth maps, instance segmentation, and class-level semantic labels across multiple virtual scenes and weather conditions. Each frame is rendered from multiple cameras and includes perfect ground truth, making it ideal for depth completion tasks. In our experiments, we use data from the front-facing camera (Camera_0), resizing all inputs to $256\times 512$ and applying an 80/20 split for training and validation 1.